OpenStudy (anonymous):

The number of blues was two more than the total number of reds and greens. Three times the number of reds added to twice the number of greens was twice the number of blues. If there were 22 total. How many were there of each color?

OpenStudy (anonymous):

b=2(r+g) 3r+2g=2b r+g+b=22

OpenStudy (anonymous):

b: blues; g: greens; r: reds b = r + g + 2 2b = 3r + 2g r + g + b = 22 First equation: b - 2 = r + g Third equation: b - 2 + b = 22 --> 2b = 24 --> b = 12 First equation: 12 = r + g + 2 --> r + g = 10 --> 3r + 3g = 30 Second equation: 24 = 3r + 2g --> 3r + 2g = 24 First equation - second equation: g = 6 6 + r = 10 --> r = 4 12 green, 4 red, 6 blue

OpenStudy (anonymous):

oh whoops mine was off